OpenStudy (lgbasallote):

LGBADERIVATIVE LEVEL 2: \[\huge y = \frac{e^x - e^{-x}}{e^x + e^{-x}}\]

OpenStudy (lgbasallote):

okay..this time idk how to do this but it looks like sinh or cosh idk

OpenStudy (zarkon):

\[y=\tanh(x)\]

OpenStudy (lgbasallote):

how?

OpenStudy (zarkon):

by definition I didn't give you the answer..just telling what your original equation is equivalent to

OpenStudy (lgbasallote):

yeah i know but how is that thingy be tanh

OpenStudy (zarkon):

\[\tanh(x)=\frac{\sinh(x)}{\cosh(x)}\]

OpenStudy (lgbasallote):

could it be that \[e^x - e^{-x} \implies \sinh (x)\] \[e^x + e^{-x} \implies cosh (x)?\]

OpenStudy (zarkon):

\[\sinh(x)=\frac{e^x-e^{-x}}{2}\]

OpenStudy (lgbasallote):

ohh hmm anyway regarding this... \[\large \frac{d}{dx} \tanh (x) = \ln |\cosh (x)|?\]

OpenStudy (zarkon):

are you integrating or differentiating?

OpenStudy (lgbasallote):

uhh oops lol

OpenStudy (lgbasallote):

so it would be \[\Large \frac{\cosh ^2 (x) - \sinh^2 (x) }{\cosh^2 (x)}\] that right??

OpenStudy (zarkon):

yes...or you can just write \(\text{sech}^2(x)\)

OpenStudy (lgbasallote):

\[\huge 1 - \tanh^2 (x) = \sec h^2 (x)?\]

OpenStudy (zarkon):

yes

OpenStudy (lgbasallote):

okay thanks! :D